# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ 4.3.1 Concentration of the stationary measure
We are interested in the behaviour of the stationary measure of KIPLD. This is a non-standard underdamped Langevin diffusion, thus, we need to first identify the stationary measure of the system. Recall that we are only interested in $\theta$-marginal of this stationary measure and we denote it by $\pi_{\Theta}$. We first have the following proposition.
@@
Under H1, the function $\theta\mapsto p_{\theta}(y)$ is $\mu$-strongly log concave. Furthermore, we have
$W_{2}(\pi_{\Theta},\delta_{\bar{\theta}_{\star}})\leq\sqrt{\frac{2d_{\theta}}{\mu N}},$
@@
See Appendix B.2. $\Box$
This shows that there is an explicit rate of convergence of the $\theta$-marginal of the stationary measure of KIPLD to the MMLE solution $\bar{\theta}_{\star}$. This is a key result, as it shows that the stationary measure of the system is concentrating on the MMLE solution as $N$ grows. We note that such results are also potentially possible under nonconvex settings [71, 1].
@@ 4.3.2 Convergence of the KIPLD to the stationary measure
We next demonstrate that the KIPLD converges to its stationary measure $\pi_{\Theta}$ exponentially fast in the strongly log-concave case. Indeed, knowing that the system KIPLD is ergodic, we would like to quantify the rate at which this mixing occurs, specifically at what rate the $W_{2}$ distance of the process and its stationary distribution converges for a possibly large range of $\gamma$.
@@
Let $(\bm{\theta}_{t})_{t\geq 0}$ be the $\theta$-marginal of a solution to the KIPLD, initialised at $\mathbf{Z}_{0}\sim\nu\leavevmode\nobreak\ \otimes\leavevmode\nobreak\ \mathcal{N}(0,\mathbb{I}_{d_{z}})$. Then under H1 and H2 and for $\gamma\geq\sqrt{\mu+L}$, we have
@@
See Appendix B.3. $\Box$
This proposition shows that, under the conditions expressed in Proposition 3, the KIPLD SDE converges exponentially fast to the target measure, which concentrates on the minimum as outlined in Proposition 2.
@@ 4.3.3 Nonasymptotic analysis of KIPLMC1
In this section we present our first main result showing the convergence rate of KIPLMC1. The result is provided in the following theorem.
@@
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq\sqrt{2}\left(1-\frac{3\mu\eta}{4\gamma}\right)^{n}{\mathbb{E}}[\|Z_{0}-\bar{Z}_{\star}\|^{2}]^{1/2}+\sqrt{2}\frac{L\eta}{\mu}\sqrt{\frac{d_{\theta}+Nd_{x}}{N}}+\sqrt{\frac{2d_{\theta}}{\mu N}},$
@@
See Appendix B.4. $\Box$
The result follows from the fact that the target measure $\pi_{\Theta}$ of our system concentrates on the maximiser which is given by Proposition 2 and using the results in [25] for kinetic Langevin diffusions and the exponential integrator discretisation. Further, we replicate the choice made in [25] of initialising the momentum particle according to $\mathcal{N}(0,\mathbb{I}_{d_{z}})$, the stationary measure under the momentum marginal.
@@ 4.3.4 Nonasymptotic analysis of KIPLMC2
In the following section, we present our second main result, showing the convergence rate of KIPLMC2. The rate is explicitly given in the subsequent theorem.
@@
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq C\left(\left(1-\frac{\eta\mu}{3\gamma}\right)^{\frac{n}{2}}\mathbb{E}[\|Z_{0}-\bar{Z}^{\star}\|^{2}]^{1/2}+\eta\sqrt{2(Nd_{x}+d_{\theta})}\frac{6\gamma K}{\mu}\right)+\sqrt{\frac{2d_{\theta}}{\mu N}},$
where $Z_{0}=(\theta_{0},N^{-1/2}X^{1}_{0},\dots,N^{-1/2}X^{N}_{0})^{\intercal}$, the initialisation step of the KIPLMC2, $\bar{Z}_{\star}\sim\tilde{\pi}$ and $\tilde{\pi}$ is the extended target measure described in Lemma A.3. The constants $C$ and $K$ are given as
$C=\sqrt{3}(\sqrt{L}\lor\sqrt{L}^{-1}),\qquad K=L\left(1+e^{L\eta^{2}}\left(\frac{\eta}{6}+\frac{\eta^{2}L}{24}\right)\right)\left(1+\frac{\eta L}{2\sqrt{\mu}}\right).$
@@
See Appendix B.5. $\Box$
This is again is a simple application of the triangle inequality for the concentration of $\pi_{\Theta}$ from Proposition 2 and the convergence results for kinetic Langevin diffusions approximated by a second-order splitting scheme from [53].
4.4 Algorithmic Complexity and Acceleration
In this section, we show that the proposed underdamped diffusion accelerates the overdamped algorithms proposed for MMLE (i.e. IPLA [3] and PGD [47]) in the same way the KLMC methods accelerate overdamped samplers such as ULA. We note that we only discuss here the impact of dimension dependence and the target accuracy $\varepsilon$.
4.4.1 Algorithmic Complexity of KIPLMC1
To provide the algorithmic complexity of KIPLMC1, it is instructive to write the bound provided in Theorem 1 free of constants that are not of interest to us:
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\lesssim\left(1-\frac{3\mu\eta}{4\gamma}\right)^{n}+\eta\sqrt{\frac{d_{\theta}+Nd_{x}}{N}}+\sqrt{\frac{d_{\theta}}{N}},$ (12)
where the notation $\lesssim$ hides the constants that are not of interest [^4] Note that we hide the dimension dependence of initial error $\mathbb{E}[\|Z_{0}-Z^{\star}\|^{2}]^{1/2}$ in this analysis. This dependence can be alleviated using a warm-start strategy in practice, e.g., running an optimiser first to initialise our algorithm, see, e.g., [14, Corollary 8].. In (12), we first set $N=\mathcal{O}(d_{\theta}\varepsilon^{-2})$ which ensures the last term in (12) is $\mathcal{O}(\varepsilon)$. Next, set $\eta=\mathcal{O}(\varepsilon d_{x}^{-1/2})$ which ensures that the second term is $\mathcal{O}(\varepsilon)$. It is then easy to see that $n=\widetilde{\mathcal{O}}(\varepsilon^{-1}d_{x}^{1/2})$ steps are enough to obtain ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq\varepsilon$.
Remark 1.
This dependency needs to be compared with IPLA and PGD. For the former, we note that our dependence to dimension for $N$ is identical to IPLA and the dependence of the number of steps is significantly improved, i.e., we require $n=\widetilde{\mathcal{O}}({d_{x}}^{1/2}\varepsilon^{-1})$ whereas IPLA requires $n=\widetilde{\mathcal{O}}({d_{x}}\varepsilon^{-2})$ steps for the same accuracy. Regarding the latter method, PGD, our dependence to number of particles is different, as PGD requires $N=\mathcal{O}(d_{x}\varepsilon^{-2})$ (but this difference is a more general difference stemming from the different analysis techniques, as IPLA also has the same dependence as us to the number of particles $N$, see, e.g., Corollary 8 and the subsequent discussion in [14]). However, in terms of the number of steps necessary to attain $\varepsilon$-error, we have the same improvement in dimension dependence in $d_{x}$ and $\varepsilon$ as we have compared to IPLA.
4.4.2 Algorithmic Complexity of KIPLMC2
Using Theorem 2, we can write for KIPLMC2 that
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\lesssim\left(1-\frac{\eta\mu}{3\gamma}\right)^{\frac{n}{2}}+\eta\sqrt{Nd_{x}+d_{\theta}}+\sqrt{\frac{d_{\theta}}{N}},$ (13)
Similarly to the analysis provided for KIPLMC1, we first set $N=\mathcal{O}(\varepsilon^{-2}d_{\theta})$ to obtain $\mathcal{O}(\varepsilon)$ dependence in the last term of (13) and $\eta=\mathcal{O}(\varepsilon^{2}({d_{x}d_{\theta}})^{-1/2})$ to obtain the same dependence in the second term of (13). Finally, setting $n=\widetilde{\mathcal{O}}(({d_{x}d_{\theta}})^{1/2}\varepsilon^{-2})$ brings us an accuracy of $\varepsilon$.
@@ 4.4.2 Algorithmic Complexity of KIPLMC2
In general, to avoid poor dependence on $N$, the error bounds of numerical discretisations must depend on $L$ and $\mu$ with the same order. This ensures that the factors cancel out appropriately, preventing any undesirable effects. This is achieved by the analysis of the exponential integrator in [25] with a linear dependence on the condition number $L/\mu$, thus our KIPLMC1 bound does not have a bad dependence to $N$. This enables us to obtain improved algorithmic complexity as discussed in Remark 1. However, for KIPLMC2, the OBABO scheme does not have this dependence on the condition number (see [53, Table 2]) under just H1 and H2. This results in a bound in Theorem 2 that grows as $N$ grows. This is reflected therefore in the algorithmic complexity as KIPLMC2 theoretically requires $n=\widetilde{\mathcal{O}}(\sqrt{d_{x}d_{\theta}}\varepsilon^{-2})$ steps to achieve $\varepsilon$ error, compared to IPLA which requires $\widetilde{\mathcal{O}}(d_{x}\varepsilon^{-2})$ steps. In this setting, the improvement (in dimension) only happens in the case where $d_{\theta}\leq d_{x}$, i.e., the dimension of parameter of the statistical model is less than the dimension of the latent variables.
step count ($n$) $N$ $\eta$
@@
Table 1: Comparison of orders required to achieve an error of order $\varepsilon$ in $W_{2}$, for step count, particle number $N$ and step size $\eta$. The IPLA case holds for $\delta>0$.
As summarised in Table 1 and Remark 2, the KIPLMC1 with a better order of convergence than the other algorithms, whilst the KIPLMC2 algorithm only has an improved rate when $d_{\theta}<d_{x}$. Indeed, the KIPLMC1 manages to fully exploit the accelerated convergence rate of the KLMC, due to the reduced sensitivity of the error on $N$. For a further discussion of a comparison of these scalings and those obtained by the Langevin samplers see [25, Sec. 3].
@@ 5 Experiments
In the following section a comparison will be made between the empirical results of KIPLMC1, KIPLMC2 algorithms, as well as, that of MPGD from [49]. We note that, in [49], the MPGD is not solely the discretisation of the KIPLD-like SDE using the exponential integrator which the authors of [49] found unstable. The MPGD proposed by [49] includes a gradient correction term. For a fair comparison, we implement both MPGD with gradient correction (which is referred to as MPGD) and MPGD without the gradient correction which we refer to as momentum particle gradient descent with no correction (MPGDnc). Our KIPLMC1 method is comparable to MPGDnc rather than MPGD which includes a gradient correction step, making theoretical analysis nontrivial.
In the following sections, we will consider three examples: (i) Bayesian logistic regression for synthetic data, (ii) Bayesian logistic regression on Wisconsin Cancer dataset, and (iii) a Bayesian Neural Network (BNN) example as a more challenging, non-convex example.
@@ 5.1 Bayesian Logistic Regression on Synthetic Data
We follow the experimental setting in [47] and [3] and start with comparisons between algorithms on a synthetic dataset for which we know the true solution. More precisely, consider the Bayesian logistic regression model
$p_{\theta}(x)=\mathcal{N}(x;\theta,\sigma^{2}I_{d_{x}}),\qquad p(y|x)=\prod_{j=1}^{d_{y}}s(v_{j}^{\intercal}x)^{y_{j}}(1-s(v_{j}^{\intercal}x))^{1-y_{j}}.$
@@
$p_{\theta}(y)=\frac{1}{(2\pi\sigma^{2})^{d_{x}/2}}\int_{\mathbb{R}^{d_{x}}}\left(\prod_{j=1}^{d_{y}}s(v_{j}^{\intercal}x)^{y_{j}}(1-s(v_{j}^{\intercal}x))^{1-y_{j}}\right)\exp\left(\frac{-\|x-\theta\|^{2}}{2\sigma^{2}}\right)\mathrm{d}x.$
@@
$\nabla_{\theta}U(\theta,x)=-\frac{x-\theta}{\sigma^{2}},\qquad\nabla_{x}U(\theta,x)=\frac{x-\theta}{\sigma^{2}}-\sum_{j=1}^{d_{y}}(y_{j}-s(v_{j}^{\intercal}x))v_{j}.$ (14)
Figure 1: Parameter estimate comparison. We compare the performance of the MPGDnc, KIPLMC1, and KIPLMC2 algorithms on the synthetic dataset with true $\bar{\theta}_{\star}\in[1,2,3]$. We observe the desired convergence of behaviours for larger values of $N$. For all the algorithms, with the chosen $\gamma$, we observe momentum effects, which extend to the noise, as can be seen in the oscillations in the low particle number regimes. In this example $d_{x}=d_{\theta}=3$, $d_{y}=500$ and $\gamma=1.2$. $Z_{0}$ is set as $\delta$, the point measure on the origin.
@@
This follows from the fact that the logistic function is Lipschitz continuous with constant ${1}/{4}$ [3]. Hence, H2 is satisfied.
@@
The sum is positive definite and the matrix was shown by [47] to have $2d_{\theta}$ positive eigenvalues and so it follows that $\nabla^{2}U$ is positive definite. Hence $U$ is strictly convex and in theory no lower bound for strong convexity constant exists. We show however this is not a problem for our practical implementations.
In Fig. 1 we can see the difference in the behaviours between the algorithms. Most notably, the KIPLMC2 algorithm exhibits comparable levels of variance as the MPGDnc algorithm in the $\theta$-dimension, whilst the KIPLMC1 algorithm has much greater variance (roughly 100 times greater). It is interesting to note that the KIPLMC2 algorithm preserves many theoretical properties of the KIPLMC1 algorithm, with empirical variance reduction properties. As $N$ grows the $\theta$ iterates concentrate onto the MMLE $\bar{\theta}_{\star}$ for all algorithms. This behaviour can be seen in more detail in Fig. 2 (b), in which we can observe the variance changing with rate $\mathcal{O}(1/N)$. Further, we make a comparison of the performance of the MPGDnc and KIPLMC2 in Fig. 2 (a) with the Area Below the Curve (ABC) metric (see C.1 for more detail). This metric is a signed, weighted difference between the performances of the two algorithms: the higher the value, the better KIPLMC2 performs compared to MPGDnc; 0 is when they perform the same. We observe that the best comparative performance for the KIPLMC2 algorithm arises in the edge cases where we have large step-size and small particle number.
Note that the momentum effects of the KIPLMC1 and KIPLMC2 algorithms has been dampened through a specific choice of $\gamma$. In training it was observed that the convergence rate is very sensitive to the choice of $\gamma$, where choices too small exhibit large momentum effects and too large lead to slow convergence. The choice of $\gamma$ here is far from optimal, but allows us to observe the strength of the proposed algorithms.
(a) Comparison of MPGDnc and KIPLMC2
(b) Variance of $\theta_{n}$ against $N$
Figure 2: Comparison over hyper-parameters. These figures compare the performance of MPGDnc with the proposed algorithms. (a) shows the Area Between the Curve (ABC) values between MPGDnc and KIPLMC2 for a variety of hyper-parameter combinations. (discussed in C.1). (b) makes a comparison over 20 Monte Carlo simulations of the algorithms’ sample variances, using the last 500 steps of each simulation. The scales are logarithmic and grid-lines corresponding to $\frac{1}{N}$ are provided to highlight the rate of convergence. Where not specified otherwise, simulations are run with $\eta=0.01$, $\gamma=0.7$ and $N=50$.

### v3

@@ 4.3.1 Concentration of the stationary measure
This is a non-standard underdamped Langevin diffusion, thus, we need to first identify the stationary measure of the system. Recall that we are only interested in $\theta$-marginal of this stationary measure and we denote it by $\pi_{\Theta}$. We first have the following proposition.
@@
Under H1, the function $\theta\mapsto p_{\theta}(y)$ is $\mu$-strongly log-concave. Furthermore, we have
$W_{2}(\pi_{\Theta},\delta_{\bar{\theta}_{\star}})\leq\sqrt{\frac{d_{\theta}}{\mu N}},$
@@
See Appendix C.2. $\Box$
We note here that this result is optimal, as it is achieved when $\pi_{\Theta}$ is Gaussian and $N\mu$ is the tight lower bound on the eigenvalues of the covariance matrix. We note that such results are also potentially possible under non-convex settings [53, 3].
@@ 4.3.2 Convergence of the KIPLD to the stationary measure
We next demonstrate that the KIPLD converges to its stationary measure $\pi_{\Theta}$ exponentially fast in the strongly log-concave case.
@@
Let $(\bm{\theta}_{t})_{t\geq 0}$ be the $\theta$-marginal of a solution to the KIPLD, initialised at $\mathbf{Z}_{0}\sim\nu\penalty 10000\ \otimes\penalty 10000\ \mathcal{N}(0,\mathbb{I}_{d_{z}})$. Then under H1 and H2 and for $\gamma\geq\sqrt{\mu+L}$, we have
@@
See Appendix C.3. $\Box$
@@ 4.3.3 Nonasymptotic analysis of KIPLMC1
In this section we present our first main result, the convergence rate of KIPLMC1.
@@
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}$ $\leq C_{1}\left(1-\frac{0.75\mu\eta}{\gamma}\right)^{n}+\eta C_{2}+\sqrt{\frac{C_{3}}{N}},$
where
$C_{1}=\sqrt{2}{\mathbb{E}}[\|Z_{0}-\bar{Z}_{\star}\|^{2}]^{1/2},\quad C_{2}=\sqrt{2}\frac{L}{\mu}\sqrt{\frac{d_{\theta}+Nd_{x}}{N}},\quad C_{3}=\frac{d_{\theta}}{\mu},$
@@
See Appendix C.4. $\Box$
To provide the algorithmic complexity of KIPLMC1, we can inspect the bound provided in Theorem 1. To obtain a complexity result, we first set $N=\mathcal{O}(d_{\theta}\varepsilon^{-2})$ which ensures the last term in the bound in Theorem 1 is $\mathcal{O}(\varepsilon)$. Next, set $\eta=\mathcal{O}(\varepsilon d_{x}^{-1/2})$ which ensures that the second term is $\mathcal{O}(\varepsilon)$. It is then easy to see that $n=\widetilde{\mathcal{O}}(\varepsilon^{-1}d_{x}^{1/2})$ steps are enough to obtain ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq\varepsilon$.
Remark 1.
This dependency needs to be compared with IPLA and PGD. For the former, we note that our dependence to dimension for $N$ is identical to IPLA and the dependence of the number of steps is significantly improved, i.e., we require $n=\widetilde{\mathcal{O}}({d_{x}}^{1/2}\varepsilon^{-1})$ whereas IPLA requires $n=\widetilde{\mathcal{O}}({d_{x}}\varepsilon^{-2})$ steps for the same accuracy. Regarding the latter method, PGD, our dependence to number of particles is different, as PGD requires $N=\mathcal{O}(d_{x}\varepsilon^{-2})$ (but this difference is a more general difference stemming from the different analysis techniques, as IPLA also has the same dependence as us to the number of particles $N$, see, e.g., Corollary 8 and the subsequent discussion in Caprio et al. [12]). However, in terms of the number of steps necessary to attain $\varepsilon$-error, we have the same improvement in dimension dependence in $d_{x}$ and $\varepsilon$ as we have compared to IPLA.
@@ 4.3.4 Nonasymptotic analysis of KIPLMC2
We present our second main result, showing the convergence rate of KIPLMC2.
@@
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq\tilde{C}_{1}\left(1-\frac{\eta\mu}{3\gamma}\right)^{n/2}+\tilde{C}_{2}\eta+\sqrt{\frac{\tilde{C}_{3}}{N}},$
where $Z_{0}=(\theta_{0},N^{-1/2}X^{1}_{0},\dots,N^{-1/2}X^{N}_{0})^{\intercal}$, the initialisation step of the KIPLMC2, $\bar{Z}_{\star}\sim\tilde{\pi}$ and $\tilde{\pi}$ is the extended target measure described in Lemma A.1. The constants $\tilde{C}_{1}$, $\tilde{C}_{2}$, $\tilde{C}_{3}$ are given as
$\tilde{C}_{1}=\sqrt{3}(\sqrt{L}\lor\sqrt{L}^{-1})\mathbb{E}[\|Z_{0}-\bar{Z}^{\star}\|^{2}]^{1/2},\quad\tilde{C}_{2}=\sqrt{2(Nd_{x}+d_{\theta})}\frac{6\gamma K}{\mu}\sqrt{3}(\sqrt{L}\lor\sqrt{L}^{-1}),\quad\tilde{C}_{3}=\frac{d_{\theta}}{\mu},$
where
$K$ $=L\left(1+e^{L\eta^{2}}\left(\frac{\eta}{6}+\frac{\eta^{2}L}{24}\right)\right)\left(1+\frac{\eta L}{2\sqrt{\mu}}\right).$
@@
See Appendix C.5. $\Box$
To obtain complexity, similarly to the analysis provided for KIPLMC1, we first set $N=\mathcal{O}(\varepsilon^{-2}d_{\theta})$ to obtain $\mathcal{O}(\varepsilon)$ dependence in the last term of the bound in Theorem 2, and $\eta=\mathcal{O}(\varepsilon^{2}({d_{x}d_{\theta}})^{-1/2})$ to obtain the same dependence in the second term of the bound in Theorem 2. Finally, setting $n=\widetilde{\mathcal{O}}(({d_{x}d_{\theta}})^{1/2}\varepsilon^{-2})$ brings us an accuracy of $\varepsilon$.
@@
In general, to avoid poor dependence on $N$, the error bounds of numerical discretisations must depend on $L$ and $\mu$ with the same order. This ensures that the factors cancel out appropriately, preventing any undesirable effects. This is achieved by the analysis of the exponential integrator in Dalalyan and Riou-Durand [19] with a linear dependence on the condition number $L/\mu$, thus our KIPLMC1 bound does not have a bad dependence to $N$. This enables us to obtain improved algorithmic complexity as discussed in Remark 1. However, for KIPLMC2, the OBABO scheme does not have this dependence on the condition number (see, e.g., Monmarché [41, Table 2]) under just H1 and H2. This results in a bound in Theorem 2 that grows as $N$ grows. This is reflected therefore in the algorithmic complexity as KIPLMC2 theoretically requires $n=\widetilde{\mathcal{O}}(\sqrt{d_{x}d_{\theta}}\varepsilon^{-2})$ steps to achieve $\varepsilon$ error, compared to IPLA which requires $\widetilde{\mathcal{O}}(d_{x}\varepsilon^{-2})$ steps. In this setting, the improvement (in dimension) only happens in the case where $d_{\theta}\leq d_{x}$, i.e., the dimension of parameter of the statistical model is less than the dimension of the latent variables.
step count ($n$) NN $\eta$
@@
Table 1: Comparison of orders required to achieve an error of order $\varepsilon$ in $W_{2}$, for step count, particle number NN and step size $\eta$.
See Table 1 for a summary of our discussion.
@@ 5 Experiments
In the following section a comparison will be made between the empirical results of KIPLMC1, KIPLMC2 algorithms, as well as, those of PGD from Kuntz et al. [35], IPLA from Akyildiz et al. [1] and MPGD from Lim et al. [37]. We note that, in Lim et al. [37], the MPGD is not solely the discretisation of the KIPLD-like SDE using the exponential integrator, but includes a gradient correction term, making theoretical analysis and a direct comparison with our methods non-trivial (for a comparison of computational cost see Sec. D.2). To this end, we will also compare these methods to MPGDnc, the MPGD algorithm implemented without gradient correction, so that the number of gradient computations is the same as KIPLMC1.
In the following sections, we will consider three examples: (i) Bayesian logistic regression on a synthetic dataset, (ii) Bayesian logistic regression on the Wisconsin Cancer dataset, and (iii) a Bayesian Neural Network (BNN) example as a more challenging, non-convex example on the MNIST dataset.
@@ 5.1 Bayesian Logistic Regression on Synthetic Data
We follow the experimental setting in Kuntz et al. [35] and Akyildiz et al. [1] and start with comparisons between algorithms on a synthetic dataset for which we know the true solution. More precisely, consider the Bayesian logistic regression model
$p_{\theta}(x)=\mathcal{N}(x;\theta,\sigma^{2}I_{d_{x}}),\quad p(y|x)=\prod_{j=1}^{d_{y}}s(v_{j}^{\intercal}x)^{y_{j}}(1-s(v_{j}^{\intercal}x))^{1-y_{j}}.$
@@
$p_{\theta}(y)=\int\left(\prod_{j=1}^{d_{y}}s(v_{j}^{\intercal}x)^{y_{j}}(1-s(v_{j}^{\intercal}x))^{1-y_{j}}\right)p_{\theta}(x)\mathrm{d}x.$
@@
$\nabla_{\theta}U(\theta,x)=-\frac{x-\theta}{\sigma^{2}},\quad\nabla_{x}U(\theta,x)=\frac{x-\theta}{\sigma^{2}}-\sum_{j=1}^{d_{y}}(y_{j}-s(v_{j}^{\intercal}x))v_{j}.$ (10)
Figure 1: Parameter estimate comparison. We compare the performance of the PGD, IPLA, MPGD, KIPLMC1, and KIPLMC2 algorithms on the synthetic dataset with true $\bar{\theta}_{\star}=\mathbbm{1}_{d_{x}}$. We observe the desired convergence of behaviours for larger values of $N$ in the parameter space, mean-square error and the posterior variance of $X_{k}^{i}$. In this example $d_{x}=d_{\theta}=3$, $d_{y}=100$ and $\gamma=2.2$. Here we choose $\theta_{0},X_{0}^{i},V^{\theta}_{0},V^{x_{i}}_{0}\sim\mathcal{N}(0,0.1)$ and run 100 Monte Carlo simulations.
@@
This follows from the fact that the logistic function is Lipschitz continuous with constant 1/4{1}/{4} Akyildiz et al. [1]. Hence, H2 is satisfied.
@@
The sum is positive definite and the matrix was shown by Kuntz et al. [35] to have $2d_{\theta}$ positive eigenvalues and so it follows that $\nabla^{2}U$ is positive definite. Hence $U$ is strictly convex and in theory no lower bound for strong convexity constant exists. We show however this is not a problem for our practical implementations.
Figure 2: Comparison over hyper-parameters. (a) shows the Area Between the Curve (ABC) values between IPLA and KIPLMC2 for a variety of hyper-parameter combinations. The larger the positive value, the greater the acceleration of KIPLMC2 with respect to IPLA (discussed in C.1). (b) makes a comparison over 100 Monte Carlo simulations of the algorithms’ sample variances, using the last 100 steps of each simulation. The grid-lines corresponding to $1/N$ are provided to highlight the rate of convergence. Where not specified otherwise, simulations are run with $M=2000$, $\eta=0.005$, $\gamma=2.2$ and $N=100$.
Figure 3: Wisconsin Dataset. The performance of PGD, IPLA, MPGD, KIPLMC1 and KIPLMC2 algorithms are compared on a logistic regression experiment for the Wisconsin Cancer Dataset. In (a), we show the behaviour of the $\theta_{n}$ iterates for a small step-size and (b) shows the behaviour $\theta_{n}$ iterates for a step-size where IPLA and PGD explode. In (c) we compare the distributions of the algorithms over different step-sizes. We set $N=100$ and $\gamma=3.2$.
In Fig. 1 we can see the difference in the behaviours between the algorithms. Most notably, the KIPLMC1 and KIPLMC2 algorithms exhibit comparable levels of acceleration as the MPGD algorithm in the $\theta$-dimension, whilst the IPLA and PGD algorithms trail further behind. Interestingly, we observe (more notably for large $N$) that the noise in the $\theta$-dynamics in KIPLMC1 and KIPLMC2, as compared to MPGD, seems to dampen some of the momentum effects when $\gamma$, the friction coefficient is sub-optimal. As $N$ grows the $\theta$ iterates concentrate onto the MMLE $\bar{\theta}_{\star}$ for all algorithms. This behaviour can be seen in more detail in Fig. 2 (b), in which we can verify the variance changing with rate $\mathcal{O}(1/N)$. Further, we make a comparison of the performance of the IPLA and KIPLMC2 in Fig. 2 (a) with the Area Below the Curve (ABC) metric (see D.3 for more detail). This metric is a signed, weighted difference between the performances of the two algorithms: the higher the value, the better KIPLMC2 performs compared to IPLA; 0 is when they perform the same. We observe that the acceleration of KIPLMC2 can be best observed in specific $\gamma$ regimes, which ensure that the momentum is close to critical dampening.
Note that the momentum effects of the KIPLMC1 and KIPLMC2 algorithms has been dampened through a specific choice of $\gamma$. In training it was observed that the convergence rate is very sensitive to the choice of $\gamma$, where choices too small, exhibit large momentum effects, and too large, lead to slow convergence (see also Lim et al. [37] for more discussion on the role of the momentum parameter for learning with KLMC). The choice of $\gamma$ here is far from optimal, but allows us to observe the strength of the proposed algorithms.